OpenStudy (trwatkins1):

Jeremy made 65% of the free throws he attempted in one basketball season. He made 39 free throws. How many free throws did Jeremy attempt?

OpenStudy (trwatkins1):

@mathstudent55

OpenStudy (trwatkins1):

@YoungStudier

OpenStudy (trwatkins1):

@truwhovian11 @Michele_Laino @Awolflover1

OpenStudy (trwatkins1):

@hartnn

hartnn (hartnn):

made = 65% of attempted

hartnn (hartnn):

39 = (65/100)*x solve for x :)

OpenStudy (trwatkins1):

so how would i do that

hartnn (hartnn):

multiply both sides by 100 divide both sides by 65

OpenStudy (trwatkins1):

ah

hartnn (hartnn):

x= 3900/65 = ...

OpenStudy (trwatkins1):

6

hartnn (hartnn):

6 or 60? :P

OpenStudy (trwatkins1):

60

OpenStudy (trwatkins1):

my 0 key is broken

hartnn (hartnn):

yup he attempted 60 throws :)

OpenStudy (trwatkins1):

An eagle was at an altitude of 1450 m. The eagle decreased its altitude to 1200 m. What was the percent decrease in the altitude? Round to the nearest percent.

hartnn (hartnn):

\(\% decrease = \dfrac{old - new}{old}\times 100 = \dfrac{1450-1200}{1450}\times 100 =...\)

OpenStudy (trwatkins1):

17

hartnn (hartnn):

good!

OpenStudy (trwatkins1):

wow im suprised i got that right XD

OpenStudy (trwatkins1):

Are x and y proportional? y = 5.7x

hartnn (hartnn):

when x = ky, we say x is directly proportional to y :)

hartnn (hartnn):

or y= kx

OpenStudy (trwatkins1):

so yes?

hartnn (hartnn):

right

OpenStudy (trwatkins1):

@hartnn when you get a sec

hartnn (hartnn):

do you see that servings = 7* number of cups always!

OpenStudy (trwatkins1):

ok

hartnn (hartnn):

its of the form, y =kx so proportional where k = 7 is the constant of proportionality

OpenStudy (trwatkins1):

University representative: This week, the University is offering an additional 20% discount on any course, on top of the 30% discount you have already received. Potential student: is that 20% off the original price or 20% off the reduced price? University representative: what's the difference? Potential student: I expect the difference is over $100. If the regular price of a course is $2000, is the student’s expectation correct?

hartnn (hartnn):

can you make a new post? if somebody else is available, he/she will help you :)

OpenStudy (trwatkins1):

:( ok